A person spent Rs 345 in buying pens and copies.if each pen cost Rs 5 and copy Rs 2.50 and if the total number of things was 108,how many of each did he buy?
Answer by rfer(16322)

p+c=108
p=108-c
--------------
5p+2.50c=345
5(108-c)+2.50c=345
540-5c+2.50c=345
-2.50c=345-540
c=-195/-2.50
c=78 copies
---------------
p=108-78=30 pens
-----------------
5p+2.5c=345
5*30+2.50*78=
150+195=345 total